what is 1+1?

(p.s i know the answer just giving you an extra question for thing your welcome!)
Answer #1

Answer- 1+1 is ofcourse 2 if we add these two numbers and use mathematical equation. But we can see just numbers and combine them and don't apply the mathematical function then we can say that 1+1 = 11.

Similarly for different numbers we can use this technique like 2+3= 23 or 3+4= 34. So, we can say that 1+1= 11.